OpenStudy (secret-ninja):

An original rectangle has a length of 14 and a width of 12. What happens to the area if the new width is double the original width? Explain your answer.

OpenStudy (anonymous):

compare \(14\times 12\) with \(14\times 2\times 12\)

OpenStudy (anonymous):

14*12=168 14*24=336 So, if the width is doubled, then the whole area is doubled.

OpenStudy (secret-ninja):

Thank you!